# WaveMamba: Spatial-Spectral Wavelet Mamba for Hyperspectral Image Classification

### v1

@@ Abstract
Hyperspectral Imaging (HSI) has proven to be a powerful tool for capturing detailed spectral and spatial information across diverse applications. Despite the advancements in Deep Learning (DL) and Transformer architectures for HSI Classification (HSIC), challenges such as computational efficiency and the need for extensive labeled data persist. This paper introduces WaveMamba, a novel approach that integrates wavelet transformation with the Spatial-Spectral Mamba architecture to enhance HSIC. WaveMamba captures both local texture patterns and global contextual relationships in an end-to-end trainable model. The Wavelet-based enhanced features are then processed through the state-space architecture to model spatial-spectral relationships and temporal dependencies. The experimental results indicate that WaveMamba surpasses existing models, achieving an accuracy improvement of 4.5% on the University of Houston dataset and a 2.0% increase on the Pavia University dataset. These findings validate its effectiveness in addressing the complex data interactions inherent in HSIs. The source code will be made public at https://github.com/mahmad00.
@@ I Introduction
Hyperspectral Imaging (HSI) has emerged as a robust technique, capturing contiguous spectral information across various wavelengths. Its applications span in diverse fields such as remote sensing [1], Earth observation, urban planning, agriculture, forestry, target/object detection , mineral exploration, environmental and climate monitoring [2, 3, 4], forensic sciences [5, 6], and even food, meat, and bakery processing [7, 8, 9, 10, 11].. HSI is noted for capturing detailed spatial and spectral information, though its high spectral resolution sensors may struggle with optimal spatial resolution in complex scenarios. The extensive spectral data in HSIs presents significant challenges and opportunities for effective classification [1].
@@
Recent advancements in Deep Learning (DL) have significantly improved HSI Classification (HSIC), with Convolutional Neural Networks (CNNs) capturing essential spatial and spectral features for accurate classification [12, 13, 14]. However, their reliance on local receptive fields and substantial labeled data highlights the need for more advanced architectures to better exploit global context [15, 16]. Meanwhile, Transformer architecture has further advanced HSIC by effectively capturing long-range dependencies and global contextual information, often surpassing traditional DL models [17, 18, 19, 20, 21, 22, 23] (Comparative Methods). Despite these improvements, Transformers face challenges related to computational efficiency and the need for extensive labeled data, limiting their practical application in HSIC [23, 24, 25, 26, 27, 28, 29] (Comparative Methods).
@@
Recently, Mamba architecture, inspired by control theory’s State Space Model, achieved linear complexity scaling with sequence length, enhancing HSIC efficiency [30]. Yao et al. [31] introduced SpectralMamba, integrating Mamba with DL techniques through a Gated Spatial-Spectral Merging (GSSM) module and a Piece-wise Sequential Scanning (PSS) strategy. Huang et al. [32] proposed the Spatial-Spectral Mamba (SSMamba) architecture, utilizing spectral-spatial token generation and multiple stacked Mamba blocks for feature fusion, though further optimization is needed. Wang et al. [33] introduced the $S^{2}$amba architecture, merging spatial and spectral features with a Spatial-Spectral Mixture Gate to enhance accuracy, but requiring further optimization for complex interactions. He et al. [34] developed the 3D-Spectral-Spatial Mamba (3DSS-Mamba) architecture, capturing global dependencies while maintaining linear complexity, though optimization is needed for high-dimensional data.
@@
To address the aforementioned issues, this paper proposes WaveMamba: a novel approach that combines the strengths of wavelet transformation and the Spatial-Spectral Mamba architecture for improved HSIC. WaveMamba introduces wavelet-enhanced spatial-spectral feature extraction within the Mamba architecture, enhancing the interaction between spectral and spatial information. This results in more accurate classification compared to traditional DL methods. By extracting wavelet-based spatial-spectral features from the HSI data and inputting them into the Mamba architecture, WaveMamba captures both local and global relationships in the data, leading to improved classification accuracy. In this paper, we made the following contributions:
@@
Figure 1: In WaveMamba, the HSI cube is divided into overlapping 3D patches centered at spatial points, each covering a $P\times P$ pixel area across all spectral bands. These patches are split into spectral and spatial tokens, enhanced using a spatial-spectral gate mechanism, and transformed with Haar wavelets into four subbands that capture different frequency components and spatial features. The subbands are concatenated into a new 3D representation, which the Mamba architecture uses to capture spatial-spectral relationships and temporal dependencies for classification.
  1. 1.
    This paper presents WaveMamba, a novel method that synergistically combines wavelet transformation with Spatial-Spectral Mamba for HSIC. By leveraging wavelet-based multi-resolution analysis, WaveMamba effectively enhances the interaction between spectral and spatial information, thereby capturing a comprehensive range of data critical for accurate HSIC.
  2. 2.
@@
    We extract spatial-spectral features from the HSI cube and process them through the Mamba architecture. Overlapping 3D patches together with wavelets improves local feature extraction, capturing fine-grained details. Additionally, a gate mechanism with fully connected dense and non-linear activation layers refines spatial and spectral tokens, enhancing feature representation.
  3. 3.
@@
    This work also demonstrates that integrating wavelet-based feature extraction with the Mamba architecture’s spatial-spectral relationship modeling achieves superior accuracy compared to traditional DL, Transformers, and Mamba methods for HSIC. Furthermore, incorporating the State-Space Model captures temporal dependencies within HSI data, while $L_{2}$ regularization in the final classification stage ensures model simplicity and generalizability, reducing overfitting and enhancing performance on unseen data.
@@ II Proposed Methodology
Spatial-spectral feature enhancement is subsequently performed which refines the HSI data features by processing the spatial $\mathbf{S}$ and spectral $\mathbf{F}$ tokens using a ”spatial and spectral gate” mechanism. This mechanism involves fully connected dense layers, non-linear activation functions, and reshaped layers. The spatial gate enhances the spatial tokens $\mathbf{S}$, and the spectral gate enhances the spectral tokens $\mathbf{F}$. This process results in improved spatial and spectral token representations for subsequent analysis.
@@
where $W_{s}$ and $W_{f}$ are learned weights and $\sigma$ is the sigmoid function.
The enhanced spatial-spectral feature map $\widetilde{\mathbf{S}}^{(l)}$ and $\widetilde{\mathbf{F}}^{(l)}$ are transformed into $\mathbf{\hat{S}}^{(l)}$ and $\mathbf{\hat{F}}^{(l)}$ i.e., $\mathbf{\hat{S}}^{(l)}$ and $\mathbf{\hat{F}}^{(l)}\in\mathcal{R}^{H\times W\times C^{*}}$ decomposed into 4 wavelet sub-channels which is down-sampled through wavelet transformation. Note that, here we used the classical Haar wavelet as expressed in [35, 36]. Concretely, the wavelet transformation is applied using a lowpass filter, denoted as $f_{l}=\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$, and a highpass filter, denoted as $f_{h}=\left(\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}}\right)$, along the rows of the input spatial and spectral features $\mathbf{\hat{S}}^{(l)}$ and $\mathbf{\hat{F}}^{(l)}$. This process results in the creation of four sub-channels, namely $\mathbf{\hat{S}}^{(l)}_{l}$, $\mathbf{\hat{S}}^{(l)}_{h}$ (Spatial) and $\mathbf{\hat{F}}^{(l)}_{l}$, $\mathbf{\hat{F}}^{(l)}_{h}$ (Spectral). Subsequently, the same lowpass filter $f_{l}$ and highpass filter $f_{h}$ are employed, this time along the columns of the derived sub-channels. This leads to the formation of eight sub-channels in total: $\mathbf{\hat{S}}^{(l)}_{ll}$, $\mathbf{\hat{S}}^{(l)}_{lh}$, $\mathbf{\hat{S}}^{(l)}_{hl}$, and $\mathbf{\hat{S}}^{(l)}_{hh}$ for spatial information and similarly $\mathbf{\hat{F}}^{(l)}_{ll}$, $\mathbf{\hat{F}}^{(l)}_{lh}$, $\mathbf{\hat{F}}^{(l)}_{hl}$, and $\mathbf{\hat{F}}^{(l)}_{hh}$ for spectral information. Each of these wavelet subbands can be viewed as a down-sampled version of the original input $\widetilde{\mathbf{S}}^{(l)}$ and $\widetilde{\mathbf{F}}^{(l)}$. They collectively preserve all the input details without any loss of information. The wavelet process is explained in Algorithm 1.
Input : $\widetilde{\mathbf{S}}^{(l)}$, $\widetilde{\mathbf{F}}^{(l)}$
1 Convert $\widetilde{\mathbf{S}}^{(l)}$ and $\widetilde{\mathbf{F}}^{(l)}$ to NumPy array $\widetilde{\mathbf{S}}^{(l)}_{np}$ and $\widetilde{\mathbf{F}}^{(l)}_{np}$;
2 Initialize empty output arrays $E_{S}$ and $E_{F}$;
3 Apply Discrete Wavelet Transform (DWT) on $\widetilde{\mathbf{S}}^{(l)}_{np}$ and $\widetilde{\mathbf{F}}^{(l)}_{np}$ using ’haar’ wavelet;
4 Extract approximation and detail coefficients $\big{(}\mathbf{\hat{S}}^{(l)}_{ll},\mathbf{\hat{S}}^{(l)}_{lh},\mathbf{\hat{S}}^{(l)}_{hl},\mathbf{\hat{S}}^{(l)}_{hh}\big{)}$ and $\big{(}\mathbf{\hat{F}}^{(l)}_{ll},\mathbf{\hat{F}}^{(l)}_{lh},\mathbf{\hat{F}}^{(l)}_{hl},\mathbf{\hat{F}}^{(l)}_{hh}\big{)}$ from DWT;
5 Combine coefficients by concatenating $E_{S}$leftarrow $\mathbf{\hat{S}}^{(l)}_{ll}$, $\mathbf{\hat{S}}^{(l)}_{lh}$, $\mathbf{\hat{S}}^{(l)}_{hl}$, $\mathbf{\hat{S}}^{(l)}_{hh}$ and $E_{F}$leftarrow $\mathbf{\hat{F}}^{(l)}_{ll}$, $\mathbf{\hat{F}}^{(l)}_{lh}$, $\mathbf{\hat{F}}^{(l)}_{hl}$, $\mathbf{\hat{F}}^{(l)}_{hh}$;
6 Reshape combined coefficients individually ($E_{S}$ and $E_{F}$) to maintain 3D shape;
7 Convert Transformed back to tensors as $E_{S}$ and $E_{F}$;
8 Concatenate $E_{S}$ and $E_{F}$ as $E$;
9 Return $E$;
Algorithm 1 Wavelet Transformation
The wavelet features are linearly transformed into the state-space model. Given a sequence of wavelet features $O=(E_{1},E_{2},E_{3},\dots,E_{T})$, the state transition is computed as:
@@ III Experimental Results and Discussion
This section presents the results from testing on the University of Houston and Pavia University HSI datasets, where the model, trained over 50 epochs with Adam optimizer and a 0.001 learning rate, effectively learned patterns using a mini-batch size of 256, Mamba block embeddings of 64, state space dimensions of 128, and a 0.1 dropout rate with $l_{2}^{2}$ regularization. The model weights were randomly initialized and fine-tuned, achieving improved performance.
@@ III-A Patch Size and Train-Validation-Test Samples
The effects of training sample percentage and patch size on model performance were evaluated to determine the optimal configuration. Table I shows that varying training percentages with a fixed $4\times 4$ patch size and testing different patch sizes reveal that smaller patches capture fine details but may overfit due to noise sensitivity. Larger patches, in contrast, capture global features, improving robustness and generalization as sample sizes increase. Optimal patch size depends on dataset size: larger patches benefit smaller datasets, while smaller patches are better for larger ones. Increasing the training set size generally enhances performance, especially with larger patches, by improving pattern learning and reducing overfitting. A balanced validation sample size is crucial for effective hyperparameter tuning, as very small or very large sets can lead to suboptimal results or excessive computation. Finally, very small test sets may not reliably estimate WaveMamba’s generalization capability.
@@
III-B Experimental Results with CNN-based Models

### v2

@@ Abstract
Hyperspectral Imaging (HSI) has proven to be a powerful tool for capturing detailed spectral and spatial information across diverse applications. Despite the advancements in Deep Learning (DL) and Transformer architectures for HSI classification, challenges such as computational efficiency and the need for extensive labeled data persist. This paper introduces WaveMamba, a novel approach that integrates wavelet transformation with the spatial-spectral Mamba architecture to enhance HSI classification. WaveMamba captures both local texture patterns and global contextual relationships in an end-to-end trainable model. The Wavelet-based enhanced features are then processed through the state-space architecture to model spatial-spectral relationships and temporal dependencies. The experimental results indicate that WaveMamba surpasses existing models, achieving an accuracy improvement of 4.5% on the University of Houston dataset and a 2.0% increase on the Pavia University dataset.
@@ I Introduction
Hyperspectral Imaging (HSI) is a robust technique for capturing extensive spectral information across a continuous range of wavelengths. This capability facilitates detailed analysis and monitoring in various applications [1]. While HSI excels in providing detailed spatial and spectral data, high spectral resolution sensors may face challenges in achieving optimal spatial resolution in complex environments. As a result, the extensive spectral data presents both challenges and opportunities for effective classification [2].
@@
Recent advances in Deep Learning (DL) have significantly improved HSI classification performance [3, 4, 5, 6]. Convolutional Neural Networks (CNNs) are particularly effective at extracting essential spatial and spectral features, enabling accurate classification [7, 8, 9]. However, their reliance on local receptive fields and substantial labeled data highlights the need for more advanced architectures to exploit global context [10, 11, 12]. The introduction of Transformer architecture has further enhanced HSI classification by effectively modeling long-range dependencies and global context, often outperforming traditional DL models [13, 14, 15]. Despite these improvements, Transformers face challenges related to computational efficiency and the requirement for extensive labeled data, which limits their practical application [16].
@@
Recently, the Mamba architecture, inspired by control theory’s State Space Model, has achieved linear complexity scaling with sequence length, thereby enhancing HSI classification efficiency [17]. Yao et al. [18] introduced SpectralMamba, which integrates Mamba with DL techniques through a gated spatial-spectral merging (GSSM) module and a piece-wise sequential scanning (PSS) strategy. Huang et al. [19] proposed the spatial-spectral Mamba (SSMamba) architecture, utilizing spectral-spatial token generation and multiple stacked Mamba blocks for feature fusion, although further optimization is required. Wang et al. [20] developed the $S^{2}$ Mamba architecture, which merges spatial and spectral features using a spatial-spectral mixture gate to enhance accuracy, but further optimization is necessary for complex interactions. He et al. [21] introduced the 3D spectral-spatial Mamba (3DSS-Mamba) architecture, capturing global dependencies while maintaining linear complexity; however, optimization is still needed for high-dimensional data.
@@
To address these challenges, this paper proposes WaveMamba, a novel approach that combines wavelet transformation with the spatial-spectral Mamba architecture for improved HSI classification. WaveMamba effectively captures both local texture patterns and global contextual relationships in an end-to-end trainable model. The wavelet-enhanced features are processed through the state-space architecture, modeling spatial-spectral relationships and temporal dependencies. This results in more accurate classification compared to traditional DL methods. By extracting wavelet-based spatial-spectral features from HSI data and integrating them into the Mamba architecture, WaveMamba captures local and global relationships, leading to improved classification accuracy. This paper makes the following contributions:
@@
Figure 1: The HSI cube is divided into overlapping 3D patches further split into spectral and spatial tokens, enhanced using a gate mechanism, and transformed with Haar wavelets into four subbands that capture different frequency components and spatial features. The subbands are concatenated into a new 3D representation, which the Mamba architecture uses to capture spatial-spectral relationships and temporal dependencies.
@@
This work introduces WaveMamba, a method that integrates wavelet transformation with spatial-spectral Mamba for HSI classification. By leveraging wavelet-based multi-resolution analysis, WaveMamba enhances the interaction between spectral and spatial information, capturing a comprehensive range of data critical for accurate classification. The use of overlapping 3D patches along with wavelets improves local feature extraction, capturing fine-grained details. Additionally, a gate mechanism with fully connected dense and non-linear activation layers refines spatial and spectral tokens, enhancing feature representation.
@@
This work demonstrates that combining wavelet-based feature extraction with the Mamba architecture’s spatial-spectral relationship modeling achieves superior accuracy compared to traditional DL, Transformer, and Mamba methods. Furthermore, incorporating the State-Space Model captures temporal dependencies within HSI data, while $L_{2}$ regularization in the final classification stage ensures model simplicity and generalizability, reducing overfitting and enhancing performance on unseen data.
@@ II Proposed Methodology
Spatial-spectral feature enhancement is subsequently performed which refines the HSI data features by processing $\mathbf{S}$ and $\mathbf{F}$ tokens using a ”spatial and spectral gate” mechanism. This mechanism involves fully connected dense layers, non-linear activation functions, and reshaped layers. The spatial gate enhances $\mathbf{S}$, and the spectral gate enhances $\mathbf{F}$. This process results in improved spatial and spectral token representations for subsequent analysis.
@@
where $W_{s}$ and $W_{f}$ are learned weights and $\sigma$ is the sigmoid function. The enhanced spatial-spectral feature map $\widetilde{\mathbf{S}}^{(l)}$ and $\widetilde{\mathbf{F}}^{(l)}$ are transformed into $\mathbf{\hat{S}}^{(l)}$ and $\mathbf{\hat{F}}^{(l)}$ i.e., $\mathbf{\hat{S}}^{(l)}$ and $\mathbf{\hat{F}}^{(l)}\in\mathcal{R}^{H\times W\times C^{*}}$ decomposed into 4 wavelet [^1] We leverage wavelets in our methodology to capture multi-scale features, improving the representation of both spatial and spectral information sub-channels which are down-sampled through wavelet transformation. Note that, here we used the classical Haar wavelet as expressed in [22, 23, 24]. Concretely, the wavelet transformation is applied using a lowpass filter, denoted as $f_{l}=\left(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right)$, and a highpass filter, denoted as $f_{h}=\left(\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}}\right)$, along the rows of the input spatial and spectral features $\mathbf{\hat{S}}^{(l)}$ and $\mathbf{\hat{F}}^{(l)}$. This process results in the creation of four sub-channels, namely $\mathbf{\hat{S}}^{(l)}_{l}$, $\mathbf{\hat{S}}^{(l)}_{h}$ (Spatial) and $\mathbf{\hat{F}}^{(l)}_{l}$, $\mathbf{\hat{F}}^{(l)}_{h}$ (Spectral). Subsequently, the same lowpass filter $f_{l}$ and highpass filter $f_{h}$ are employed, this time along the columns of the derived sub-channels. This leads to the formation of eight sub-channels in total: $\mathbf{\hat{S}}^{(l)}_{ll}$, $\mathbf{\hat{S}}^{(l)}_{lh}$, $\mathbf{\hat{S}}^{(l)}_{hl}$, and $\mathbf{\hat{S}}^{(l)}_{hh}$ for spatial information and similarly $\mathbf{\hat{F}}^{(l)}_{ll}$, $\mathbf{\hat{F}}^{(l)}_{lh}$, $\mathbf{\hat{F}}^{(l)}_{hl}$, and $\mathbf{\hat{F}}^{(l)}_{hh}$ for spectral information. Each of these wavelet subbands can be viewed as a down-sampled version of the original input $\widetilde{\mathbf{S}}^{(l)}$ and $\widetilde{\mathbf{F}}^{(l)}$. They collectively preserve all the input details without any loss of information.
The wavelet features are linearly transformed into the state-space model [^2] State-space models represent systems through a set of hidden states that evolve over time. In our model, state-space modeling helps capture spatial-spectral dependencies.. Given a sequence of wavelet features $O=(E_{1},E_{2},E_{3},\dots,E_{T})$, the state transition is computed as:
@@ III Experimental Results and Discussion
For all the following results, WaveMamba trained over 50 epochs with Adam optimizer and a 0.001 learning rate, effectively learned patterns using a mini-batch size of 256, Mamba block embeddings of 64, state space dimensions of 128, and a 0.1 dropout rate with $l_{2}^{2}$ regularization. The model weights were randomly initialized and fine-tuned, achieving improved performance.
@@ III-A Patch Size and Train-Validation-Test Samples
The effects of training sample percentage and patch size on WaveMamba performance were evaluated to determine the optimal configuration. Table I shows that varying training percentages with a fixed $4\times 4$ patch size and testing different patch sizes reveal that smaller patches capture fine details but may overfit due to noise sensitivity. Larger patches, in contrast, capture global features, improving robustness and generalization as sample sizes increase. Optimal patch size depends on dataset size: larger patches benefit smaller datasets, while smaller patches are better for larger ones. Increasing the training set size generally enhances performance, especially with larger patches, by improving pattern learning and reducing overfitting. A balanced validation sample size is crucial for effective hyperparameter tuning, as very small or very large sets can lead to suboptimal results or excessive computation. Finally, very small test sets may not reliably estimate WaveMamba’s generalization capability.
@@
III-B Impact of Wavelet Transformation
We examined the impact of the wavelet transformation by comparing the model’s performance with and without wavelets, highlighting its contribution to feature enhancement as shown in Table II which provides a more detailed analysis of how the wavelets module’s absence affects overall performance, specifically focusing on performance degradation when key components like the wavelet transformation are removed. This will help emphasize the importance of each component in the model’s architecture.
TABLE II: Performance with and without Wavelet.
Class University of Houston Class Pavia University
Without With Without With
Healthy grass 94.4089 99.0415 Asphalt 96.6224 98.5223
Stressed grass 99.8405 99.5215 Meadows 99.3457 99.8712
Synthetic grass 99.1379 100 Gravel 88.9418 92.2783
Trees 96.6237 98.7138 Trees 97.4543 98.5639
Soil 98.5507 99.8389 Painted 100 99.1084
Water 93.8271 99.3827 Soil 97.4940 99.4431
Residential 92.4290 93.6908 Bitumen 97.8947 98.1954
Commercial 92.4437 95.8199 Bricks 92.3411 95.2743
Road 93.6102 96.0063 Shadows 95.7805 98.1012
Highway 96.5798 98.8599 OA 96.2083 98.6347
Railway 97.0873 98.0582 AA 97.3536 97.7065
Parking Lot 1 94.1653 98.8654 kappa 96.4942 98.1901
Parking Lot 2 64.9572 88.0341
Tennis Court 98.5981 99.5327
Running Track 99.0909 100
OA 94.9833 97.8043
AA 94.5737 97.6910
kappa 94.0900 97.6260
TABLE III: Per Class classification accuracy (Overall (OA) and Average Accuracy (AA) %) and Kappa ($\kappa$) measure for University of Houston. (WM = WaveMamba).
Class [13] [25] [26] [16] [27] [15] [28] [29] [30] [31] [32] [33] [34] [35] [36] WM
1 93.31 97.26 93.39 97.24 83.00 88.13 81.23 80.51 79.47 93.44 93.48 95.36 92.67 90.90 92.88 99.04
2 97.81 97.29 99.54 97.69 83.74 90.04 87.59 67.45 79.08 95.37 95.10 98.58 97.57 95.11 95.99 99.52
4 100.0 95.78 98.09 97.85 92.57 89.49 77.05 97.39 98.51 97.69 97.13 97.82 96.58 92.27 99.17 98.71
5 98.16 98.41 97.70 99.91 99.81 98.20 90.09 96.64 100.0 98.61 97.66 99.99 99.78 98.58 98.29 99.83
6 100.0 91.36 100.0 92.46 100.0 99.30 93.94 84.66 87.53 95.08 97.37 98.42 97.47 96.42 96.39 99.38
7 87.83 94.60 90.96 90.70 87.59 98.97 87.53 78.78 77.00 91.46 91.93 84.32 86.43 89.18 91.60 93.69
8 85.91 91.82 89.18 96.78 73.22 98.01 93.02 45.49 49.77 78.02 94.88 80.80 83.25 83.90 83.33 95.81
9 75.33 92.39 90.62 94.85 81.21 95.94 86.12 76.05 72.19 92.26 88.57 78.82 82.61 86.58 92.05 96.01
10 82.52 90.61 93.22 99.27 68.15 98.65 78.68 36.93 94.51 97.61 89.76 95.10 96.50 99.09 98.05 98.85
11 79.19 89.09 87.91 96.13 89.85 87.38 81.65 64.05 85.59 93.82 95.44 95.82 93.46 92.34 92.48 98.05
12 72.76 94.18 83.15 99.18 89.15 94.33 89.71 70.01 59.63 92.45 93.15 89.65 89.55 91.27 91.24 98.86
13 79.49 82.51 84.11 65.16 92.28 98.95 85.49 98.81 80.06 95.86 82.75 97.75 92.98 91.22 92.87 88.03
OA 88.45 93.06 93.09 96.08 86.30 93.21 86.50 75.65 82.14 93.92 93.67 92.88 92.85 92.73 94.30 97.80
AA 87.81 86.61 92.06 94.93 88.68 93.22 85.25 79.34 84.19 94.65 94.03 94.10 93.79 93.51 94.96 97.69
$\kappa$ 87.50 92.50 92.53 95.76 85.14 92.63 85.85 73.73 80.70 93.43 93.16 92.30 92.27 92.14 93.84 97.62
TABLE IV: Per Class classification accuracy (Overall (OA) and Average Accuracy (AA) %) and Kappa ($\kappa$) measure for Pavia University. (WM = WaveMamba).
Class [13] [25] [26] [16] [27] [15] [28] [29] [30] [31] [32] [33] [34] [35] [36] WM
1 92.67 95.21 95.25 96.34 90.44 94.75 98.76 78.09 76.94 97.06 98.74 86.16 87.16 90.59 95.70 98.52
2 92.79 92.54 96.64 99.51 96.63 95.54 95.46 84.67 93.70 92.29 99.51 9 4.50 95.05 94.62 94.05 99.87
3 90.60 91.18 81.74 82.74 93.50 99.83 87.75 53.47 81.10 99.97 90.81 93.94 93.95 94.39 99.61 92.27
4 98.15 97.21 95.50 96.33 90.94 93.99 91.37 75.09 78.42 97.12 92.74 88.97 92.76 84.61 98.92 98.56
5 98.28 100.0 99.59 100.0 98.71 95.82 92.13 99.83 100.0 100.0 99.53 98.95 98.85 99.18 100.0 99.11
6 93.29 99.93 93.43 92.31 98.82 97.32 97.10 52.40 64.24 99.40 90.96 96.07 99.19 99.53 99.19 99.44
7 83.01 95.75 89.24 92.06 98.13 96.58 96.47 78.74 96.27 100.0 93.80 99.58 99.08 99.34 99.93 98.19
8 84.50 97.59 88.81 90.67 93.95 95.37 86.25 84.35 58.58 98.99 89.83 86.63 91.74 97.10 98.50 95.27
9 99.77 99.47 99.53 93.30 95.24 92.18 89.27 99.60 98.16 99.47 96.82 95.54 95.49 93.95 99.96 98.10
OA 92.30 91.35 94.48 96.05 95.25 95.05 94.64 78.23 83.28 95.79 95.87 92.61 94.06 94.33 96.40 98.63
AA 88.86 85.07 93.15 93.71 95.15 93.01 92.78 78.47 83.06 98.25 94.75 93.37 94.81 94.81 98.43 97.70
$\kappa$ 89.66 88.94 92.67 94.74 93.55 94.70 93.54 71.55 77.74 94.54 94.58 90.29 92.22 92.60 95.31 98.19
III-C Experimental Results with CNN-based Models